OpenStudy (anonymous):

a line goes trough the points: T(0.0) = top P(1,2) how do I find the function of this line( its a quadratic function)?

OpenStudy (mathmate):

It is a parabola with the vertex at (0,0) and passes through P(1,2). The standard form of the parabola is y=a(x-h)^2+k where (h,k) is the vertex, and a is a constant to be determined (by the point P).

OpenStudy (anonymous):

i'll try to solve it now

OpenStudy (anonymous):

@mathmate approach is correct

OpenStudy (anonymous):

im stuck, idon't know where to start

OpenStudy (anonymous):

ok, i get it now, is the answer 2x^2?

OpenStudy (anonymous):

u asked for line equation pasiing thru 2 points?

OpenStudy (anonymous):

yes. \[y = 2x^2\], is the anxwer right? a quadratic equation

OpenStudy (anonymous):

well done

ganeshie8 (ganeshie8):

alternative solution let the quadratc be y = ax^2 + bx + c --------- (1) since its going thru T(0,0) : 0 = a(0)^2 + b(0) + c => c = 0 ------------(2) since it has top at T(0, 0) : dy/dx at x=0 is 0 => 2ax + b = 0 => 2a(0) + b = 0 => b = 0 --------------------(3) combining (1), (2) & (3) y = ax^2 ----------------(4) since its also going thru P(1, 2) 2 = a(1)^2 => 2 = a ----------------(5) plugin a=2 in (4) y = 2x^2

OpenStudy (anonymous):

thank you, i had to give the function in the form y= ax^2+bx+c, so this just what i needed

ganeshie8 (ganeshie8):

np..